# An Information-Theoretic Framework for Receiver Quantization in Communication

### v1

@@ III Achievable Rate Under Receiver Quantization: Exact and Asymptotic Results
$\mathrm{E}\left[X\overline{Y}\right]$ $=2\mathrm{E}\left[\mathrm{Re}(X)\cdot\mathrm{Re}(Y)+\mathrm{j}\cdot\mathrm{Im}(X)\cdot\mathrm{Re}(Y)\right]$ (24a)
@@
$=\frac{2\overline{h}\sigma_{x}^{2}}{\sqrt{\pi(|h|^{2}\sigma_{x}^{2}+\sigma^{2})}}\mathcal{A}$ (24c)
@@
We note that, if the impact of quantization is omitted, then the model (4) reduces to a linear Gaussian channel $Y_{n}=hX_{n}(m)+Z_{n}$, and the achievable rate of the considered transceiver architecture is $\log(1+\mathsf{SNR})$, which is also the channel capacity $C$ when the power of the input is $\sigma_{x}^{2}$. The expression of $I_{\text{GMI}}$ in Theorem 2 shows explicitly that the rate loss due to quantization is
@@
which has the same expression as $C$ except for an SNR reduction of at least $4.4$ dB (see (28) given later). For a given SNR, the reduction is determined by the parameter $\gamma$, which reflects the impact of nonlinearity and does not depend on the SNR.
Remark: Dithering, an intentional randomization technique in source quantization [41], may also be beneficial in receiver quantization; see, e.g., [31], which showed that nonsubtractive Gaussian dithering improves GMI under certain circumstances. [^5] In nonsubtractive dithering, one adds a dither signal $W_{\mathrm{d}}$ independent of the quantizer input $V$ before quantization, yielding an output $q(V+W_{\mathrm{d}})$. Subtractive dithering has an additional step that subtracts the dither signal after quantization and finally obtains $q(V+W_{\mathrm{d}})-W_{\mathrm{d}}$. See [41] for more discussion. However, in our setting, nonsubtractive Gaussian dithering is always harmful, because it is equivalent to reducing the SNR, while the GMI is a monotonically increasing function of the $\mathsf{SNR}$ since its derivative (in nats) satisfies
$\frac{\mathrm{d}I_{\textrm{GMI}}}{\mathrm{d}\mathsf{SNR}}=\frac{1-\gamma}{(1+\mathsf{SNR})(1+\gamma\mathsf{SNR})}>0.$ (27)
From Theorem 2, we immediately obtain the following corollary, which shows that $\gamma$ dominates the asymptotic behavior of performance, especially the low-SNR slope of achievable rate, and the high-SNR saturation rate.
@@
  • The parameter $\gamma$ satisfies
    $0<\gamma\leq 1-\frac{2}{\pi},$ (28)
    where the equality holds when $q(\cdot)$ is a one-bit quantizer, corresponding to an achievable rate
    $I_{\textrm{GMI}}^{\textrm{1-bit}}=\log\frac{1+\mathsf{SNR}}{1+\frac{\pi-2}{\pi}\mathsf{SNR}},$ (29)
    which converges to $\log_{2}\frac{\pi}{\pi-2}=1.4604$ bits/c.u. at high SNR (see also footnote 6). As the quantization becomes increasingly fine, we have $\gamma\to 0$ (from above), $\mathsf{SNR}_{\mathrm{e}}\to\mathsf{SNR}$ (from below), and
[…]
@@
    So the saturation rate is given by
@@
[…]
An alternative expression of (19) is given in terms of effective SNR (or signal to noise-and-distortion ratio) as
$I_{\text{GMI}}=\log\left(1+\mathsf{SNR}_{\mathrm{e}}\right),$ (34)
@@
From this expression, we see that the effect of quantization is to transfer a fraction $\gamma$ of the power to the denominator of the effective SNR. At low SNR, the power loss can be evaluated by $\mathsf{SNR}/\mathsf{SNR}_{\mathrm{e}}$, which asymptotically converges to $21/(1-\gamma)\leq\pi/$ 2, where equality holds when one-bit quantization is used. [^6] The channel capacity under one-bit receiver quantization, given by $1-H_{2}(Q(\mathsf{SNR}))$ [6, 8], has the same low-SNR asymptotic behavior. Note that the GMI is derived under Gaussian input and nearest neighbor decoder, and the capacity is achieved by antipodal signaling (the decoding rule is unrestricted). The difference becomes evident at high-SNR, where the capacity converges to $2$ bits/c.u., while the GMI converges to $1.4604$ bits/c.u.
From Theorem 2, it is simple to check that, for one-bit quantization, the gain control factor $g$ has no impact on the achievable rate. In general, the achievable rate has the following property.
$\lim\limits_{g\to 0}I_{\text{GMI}}(g)=\lim\limits_{g\to\infty}I_{\text{GMI}}(g)=I_{\textrm{GMI}}^{\textrm{1-bit}}.$ (36)
@@
Remark: We note that $I_{\textrm{GMI}}^{\textrm{1-bit}}$ is not the infimum of $I_{\textrm{GMI}}(g)$ in general. An example will be given in Sec. IV-D, Fig. 17(f).
Intuitively, a smaller MSE is preferred in quantization. However, for communications the main performance metric is the achievable rate. Thus the relationship of MSE minimization and achievable rate maximization is of interest.
Due to symmetry, we may define the normalized MSE of the quantizer $q(\cdot)$ in the channel (4) in terms of the in-phase component of $Y_{n}$ as
@@
The different expressions of $\mathsf{mse}$ and $\gamma$ show that, in general, the gain control factor that minimizes the MSE does not necessarily maximizes the achievable rate. However, for uniform quantization, the next section will show that the two optimization problems are consistent; i.e., there is a unique gain control factor (and correspondingly, a unique loading factor) that solves both problems simultaneously.
@@ IV-A Achievable Rate Evaluation under Uniform Quantization
Proposition 6: If $q(\cdot)$ is the uniform quantizer described in Sec. II-A2, then
@@
Specifically, for a uniform quantizer with a given resolution, (44) and (45) show that $\gamma$ is determined solely by the step size $\ell$ or equivalently by the loading factor $L$, which can be optimized by adjusting the gain-control factor $g$ in (4) according to the channel gain $h$.
Remark: The quantities $\mathcal{A}$ and $\mathcal{B}$ are important in our analysis. For uniform quantization, we note that the summation $\sum_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}$ in $\mathcal{A}$ is exactly the left Riemann sum of $\exp\frac{-t^{2}}{2}$ over $[0,K\ell]$ with a regular partition, and similarly, the summation $\pi\sum_{k=0}^{K-1}2k\ell^{2}Q(k\ell)$ in $\mathcal{B}$ is exactly the left Riemann sum of $2\pi tQ(t)$ over $[0,K\ell]$ with a regular partition. Therefore, for a fixed loading factor $L$ we have the following high-resolution limits:
@@
$\lim\limits_{K\to\infty}\mathcal{B}$ $=2\pi\lim\limits_{K\to\infty}\left(\sum\limits_{k=0}^{K-1}\frac{L}{K}\frac{kL}{K}Q\left(\frac{kL}{K}\right)+\frac{L^{2}}{16K^{2}}\right)$ (47a)
@@
Figure 3: Convergence of achievable rate to channel capacity.
@@
$\mathsf{mse}_{\textrm{o}}:=\int_{L}^{\infty}(t-L+\ell/2)^{2}\phi(t)\mathrm{d}t,$ (49)
@@
$\mathsf{mse}_{\textrm{g}}:=\sum\limits_{k=1}^{K}\int_{(k-1)\ell}^{k\ell}\left(t-\left(k-\frac{1}{2}\right)\ell\right)^{2}\phi(t)\mathrm{d}t.$ (50)
@@
The numerical results in Fig. 4 and Fig. 5 reveal the increasing importance of gain control (realized by an AGC module in practical systems) as the resolution decreases: 1) Under high-resolution output quantization, we only require a rough estimate of the channel gain to guarantee that the loading factor to be no less than a predefined threshold, say $4$ (from the four-sigma rule of thumb [44, 47]), so that we can stay away from the waterfall; 2) Under low-resolution output quantization, such a simple strategy may increase rate loss considerably, but perfect gain control needs accurate channel estimation which is also challenging in this case.
@@
Figure 5: Impact of loading factor on achievable rate: Fixed SNR, varying resolution.
@@ IV-B Consistency Between Rate Maximization and MSE Minimization and a GMI-MMSE Formula
From source quantization theory [46, 50] we know that, for Gaussian input, the MSE of the uniform quantizer is a strictly convex function of the step size (or loading factor) and the minimum is located at the unique solution of $\frac{\mathrm{d}\mathsf{MSE}}{\mathrm{d}\ell}=0$, denoted by $\ell^{*}$. In fact, that solution is also the unique step size that maximize the GMI; that is, for uniform quantization at the receiver there is a consistency between achievable rate maximization and MSE minimization. This consistency further leads to a GMI-MMSE formula which connects the maximum achievable rate and the MMSE of quantization in a simple and closed form. We first give the following lemma, and then use it to establish the aforementioned findings in Theorem 8.
@@
    $0<\mathsf{mmse}=\gamma(L^{*})\leq\gamma\leq 1-\frac{2}{\pi},$ (54)
@@
First, in [References, Sec. V-A], it has been proved that the MSE is minimized if and only if the loading factor is set to be the unique solution of $\frac{\mathrm{d}\mathsf{mse}}{\mathrm{d}L}=0$, denoted by $L^{*}$. The uniqueness of $L^{*}$ is confirmed by showing that the second derivative of the MSE is strictly positive. Lemma 7 implies that the loading factor that satisfies $\frac{\mathrm{d}I_{\text{GMI}}}{\mathrm{d}L}=\frac{\mathrm{d}I_{\text{GMI}}}{K\mathrm{d}\ell}=0$ is also unique. Noting that the achievable rate $I_{\text{GMI}}$ is a continuous and differentiable function of $L$, from Proposition 4 we can infer that $L^{*}$ also maximizes $I_{\text{GMI}}(L)$ (it is not difficult to exclude the other possible case that $L^{*}$ minimizes $I_{\text{GMI}}(L)$). Then the first part of Theorem 8 is proved.
@@
From the proof of Theorem 8, we can infer that $I_{\text{GMI}}(L)>I_{\textrm{GMI}}^{\textrm{1-bit}}$ holds for $0<L<\infty$. Combining this with Proposition 4 yields the following corollary.
Corollary 9: For uniform receiver quantization, the achievable rate given in (19), as a function of the loading factor $L$, satisfies
@@
where $I_{\textrm{GMI}}^{\textrm{1-bit}}$ is given in (29).
In Theorem 8 we exclude the case $b=1$. In fact, for symmetric one-bit quantization, we have $\mathcal{A}=\ell/2$, $\mathcal{B}=\pi\ell^{2}/8$, and $\gamma\equiv 1-2/\pi$ for $0<\ell<\infty$, so that gain control is unnecessary. But when
$\ell=\frac{4}{\sqrt{2\pi}}=1.5958,$ (58)
the normalized MSE achieves its minimum as $\mathsf{mmse}^{\textrm{1-bit}}=1-2/\pi$, corresponding to the upper bound in (54).
@@
Theorem 8 enables us to utilize known results on uniform quantization for minimizing the MSE. The following result follows immediately from the equivalence (53) and existing results in [67, 46]. In particular, it is shown that the optimal loading factor $L^{*}$ grows with the resolution $b$ like $2\sqrt{b\ln 2}$.
@@
For (61) and (62), we also provide new proofs of them, respectively, by asymptotic results on the achievable rate; see Sec. VI-A.
Remark: In quantization theory, the “SNR” for a quantizer is often defined as the ratio between the variance of the quantizer input and the MSE. For the optimal uniform quantizer we denote
@@
Theorem 8 implies that the maximum saturation rate can be expressed as
@@
$\mathsf{SNR}_{\textrm{e}}=\frac{1-\mathsf{mmse}}{\mathsf{mmse}\cdot\mathsf{SNR}+1}\mathsf{SNR}=\frac{(1-\mathsf{mmse})\mathcal{E}_{x}}{\mathsf{mmse}\cdot\mathcal{E}_{x}+\sigma^{2}}.$ (68)
@@
$1$ 2 1.5958 1.5958 0.3634 0.2310 4.40 1.4604 2.11
$2$ 4 1.9914 0.9957 0.1188 0.1155 9.25 3.0728 3.11
$3$ 8 2.3441 0.5860 3.7440×$\times 10^{-2}$ 4.3322×$\times 10^{-2}$ 14.27 4.7393 4.53
$4$ 16 2.6816 0.3352 1.1543×$\times 10^{-2}$ 1.4441×$\times 10^{-2}$ 19.38 6.4369 6.11
$5$ 32 3.0102 0.1881 3.4952×$\times 10^{-3}$ 4.5127×$\times 10^{-3}$ 24.57 8.1604 7.79
$6$ 64 3.3300 0.1041 1.0400×$\times 10^{-3}$ 1.3538×$\times 10^{-3}$ 29.83 9.9091 9.53
$7$ 128 3.6395 5.6868×$\times 10^{-2}$ 3.0433×$\times 10^{-4}$ 3.9486×$\times 10^{-4}$ 35.17 11.6821 11.31
$8$ 256 3.9376 3.0762×$\times 10^{-2}$ 8.7686×$\times 10^{-5}$ 1.1282×$\times 10^{-4}$ 40.57 13.4773 13.11
$9$ 512 4.2237 1.6499×$\times 10^{-2}$ 2.4919×$\times 10^{-5}$ 3.1730×$\times 10^{-5}$ 46.03 15.2924 14.94
$10$ 1024 4.4982 8.7855×$\times 10^{-3}$ 6.9970×$\times 10^{-6}$ 8.8138×$\times 10^{-6}$ 51.55 17.1248 16.79
$11$ 2048 4.7614 4.6498×$\times 10^{-3}$ 1.9444×$\times 10^{-6}$ 2.4238×$\times 10^{-6}$ 57.11 18.9722 18.65
$12$ 4096 5.0143 2.4484×$\times 10^{-3}$ 5.3554×$\times 10^{-7}$ 6.6104×$\times 10^{-7}$ 62.71 20.8325 20.53
@@
  • SNR-limited (high resolution) scenario: $b>2\log_{10}\mathsf{SNR}$ $+b_{1}$ bits and $\mathsf{SNR}_{q}\gg\mathsf{SNR}$, where $b_{1}$ can be $2$~$3$. The irreducible loss (59) is negligible, and significant rate improvement can be obtained by increasing the SNR. In fact, we have
@@
IV-C Geometry of Optimal Uniform Quantization at Receiver
We have shown that, in the standard transceiver architecture shown in Fig. 1, if the gain control factor $g$ is set appropriately so that the loading factor of the quantizer $q(\cdot)$ is equal to the optimal value $L^{*}$, then the achievable rate attains its maximum given by the GMI-MMSE formula (55). In fact, a geometrical interpretation for such optimal uniform quantization can be established. For simplification we first introduce an equivalent model shown in Fig. 7, where we let $\mathsf{X}=hX$, $\mathsf{Y}=\sigma_{v}Y=Y\cdot\sqrt{(|h|^{2}\sigma_{x}^{2}+\sigma^{2})/2}$, $\mathsf{V}=V^{\mathrm{R}}+\mathrm{j}V^{\mathrm{I}}$, $\mathsf{Z}=Z$, and $\mathsf{W}=\mathsf{Y}-\mathsf{V}$ (we write $\mathsf{W}(\mathsf{V})$ in Fig. 7 to emphasize that $\mathsf{W}$ is a function of $\mathsf{V}$), so that

### v2

@@ III Achievable Rate Under Receiver Quantization: Exact and Asymptotic Results
[…]
@@
$=\frac{2}{\sqrt{\pi}}\frac{\overline{h}\sigma_{x}^{2}}{\sqrt{|h|^{2}\sigma_{x}^{2}+\sigma^{2}}}\mathcal{A}$ (26c)
@@
If the impact of quantization is omitted, then the model (4) reduces to a linear Gaussian channel $Y_{n}=hX_{n}(m)+Z_{n}$, and the achievable rate of the considered transceiver architecture is $\log(1+\mathsf{SNR})$, which is also the channel capacity $C$ when the power of the input is $\sigma_{x}^{2}$. The expression of $I_{\text{GMI}}$ in Theorem 2 shows explicitly that the rate loss due to quantization is
@@
which has the same expression as $C$ except for a pre-SNR factor $\gamma$. For a given SNR, the rate loss is determined by $\gamma$, which reflects the impact of nonlinearity and does not depend on the SNR. For the GMI (21) given in Theorem 2, the effective SNR is
From this expression, we see that the effect of quantization is to transfer a fraction $\gamma$ of the power to the denominator of the effective SNR.
From Theorem 2, we immediately obtain the following corollary, which shows that the parameter $\gamma$ dominates the asymptotic behavior of performance, especially the low-SNR slope of achievable rate and the saturation rate.
@@
  • The parameter $\gamma$ in (21) satisfies $0<\gamma<1$. As the quantization becomes increasingly fine, we have $\gamma\to 0$ (from above), $\mathsf{SNR}_{\mathrm{e}}\to\mathsf{SNR}$ (from below), and
    $C-I_{\text{GMI}}=\mathsf{SNR}\cdot\gamma-\frac{\mathsf{SNR}^{2}}{2}\gamma^{2}+o(\gamma^{2})\;\;\textrm{nats/c.u.}$ (30)
@@
    So the saturation rate (i.e. the high-SNR limit of achievable rate) is given by
@@
    $I_{\textrm{GMI}}=(1-\gamma)\mathsf{SNR}$ $-\frac{1-\gamma^{2}}{2}\mathsf{SNR}^{2}+o\left(\mathsf{SNR}^{2}\right)\;\;\textrm{nats/c.u.}$ (33)
Proof:
The upper bound $\gamma<1$ follows from the fact that $\mathcal{A}$ and $\mathcal{B}$ are always positive (guaranteed by the setting of the quantizer in Sec. II-A-2). The lower bound $\gamma>0$ follows from the fact $\mathcal{A}^{2}<\mathcal{B}$, which can be proved by combining (26c) and (27) and then applying the Cauchy-Schwartz inequality. The upper bound is consistent with the nonnegativity of the GMI (21), while the lower bound can also be seen directly from the fact that the GMI (21) cannot exceed the capacity $C$. For increasingly fine quantization, a formal proof of $\gamma\to 0$ can be found in [References, Sec. V] (although it considered a real-valued channel, the proof goes through verbatim in our setting except for some different notations). Since $\gamma$ is always nonnegative, we can infer that $\gamma$ tends to zero from above, and thus $\mathsf{SNR}_{\mathrm{e}}$ tends to $\mathsf{SNR}$ from below. The properties (30), (31), and (33) can be obtained from the expression (21) by standard asymptotic analysis. ∎
The following result considers extreme cases of GMI with gain control.
$\lim\limits_{g\to 0}I_{\text{GMI}}(g)=\lim\limits_{g\to\infty}I_{\text{GMI}}(g)=I_{\textrm{GMI}}^{\textrm{1-bit}},$ (34)
@@
$I_{\textrm{GMI}}^{\textrm{1-bit}}=\log\frac{1+\mathsf{SNR}}{1+\frac{\pi-2}{\pi}\mathsf{SNR}}$ (35)
is the GMI under one-bit quantization.
@@
Remark: At high SNR the GMI $I_{\textrm{GMI}}^{\textrm{1-bit}}$ converges to $\log_{2}\frac{\pi}{\pi-2}=1.4604$ bits/c.u., while in the low-SNR limit it has a slope of $2/\pi$. [^6] The channel capacity under one-bit receiver quantization, given by $1-H_{2}(Q(\mathsf{SNR}))$ [6, 8], has the same low-SNR asymptotic behavior, i.e., the well-known “2-dB loss” result (see Sec. I-A). Note that the GMI is derived under Gaussian input and nearest neighbor decoder, and the capacity is achieved by antipodal signaling (the decoding rule is unrestricted). The difference becomes evident at high-SNR, where the capacity converges to $2$ bits/c.u., while the GMI converges to $1.4604$ bits/c.u. We note that $I_{\textrm{GMI}}^{\textrm{1-bit}}$ is not the infimum of $I_{\textrm{GMI}}(g)$ in general; see an example given in Sec. IV-D, Fig. 17(f).
Intuitively, a smaller MSE is preferred in quantization. However, for communications the main performance metric is the achievable rate. Thus the relationship of MSE minimization and achievable rate maximization is of interest. Due to symmetry, we may define the normalized MSE of the quantizer $q(\cdot)$ in the channel (4) in terms of the in-phase component of $Y_{n}$ as
@@
The different expressions of $\mathsf{mse}$ in (39) and $\gamma$ in (22) show that, in general, the gain control factor that minimizes the MSE does not necessarily maximize the achievable rate. However, for uniform quantization, the next section will show that the two optimization problems are consistent; i.e., there is a unique gain control factor (and correspondingly, a unique loading factor) that solves both problems simultaneously.
@@ IV-A Achievable Rate Evaluation under Uniform Quantization
Corollary 6: If $q(\cdot)$ is the uniform quantizer described in Sec. II-A2, then
@@
For a given resolution, (43) and (44) show that $\gamma$ is determined solely by the step size $\ell$ or equivalently by the loading factor $L$, which can be optimized by adjusting the gain-control factor $g$ in (4) according to the channel gain $h$.
Remark: The quantities $\mathcal{A}$ and $\mathcal{B}$ are important in our analysis. For uniform quantization, we note that the summation $\sum_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}$ in $\mathcal{A}$ is exactly the left Riemann sum of $\exp\frac{-t^{2}}{2}$ over $0,K\ell]$ with a regular partition, and similarly, the summation $\pi\sum_{k=0}^{K-1}2k\ell^{2}Q(k\ell)$ in $\mathcal{B}$ is exactly the left Riemann sum of $2\pi tQ(t)$ over $0,K\ell]$ with a regular partition. Therefore, for a fixed loading factor $L$ we have the following high-resolution limits:
@@
[…]
@@
$\mathsf{mse}_{\textrm{o}}$ $:=\int_{L}^{\infty}\left(t-L+\frac{\ell}{2}\right)^{2}\phi(t)\mathrm{d}t$ (48a)
$=\int_{K\ell}^{\infty}\left(t-\left(K-\frac{1}{2}\right)\ell\right)^{2}\phi(t)\mathrm{d}t,$ (48b)
@@
$\mathsf{mse}_{\textrm{g}}$ $:=\sum\limits_{k=1}^{K}\int_{(k-1)\ell}^{k\ell}\left(t-\left(k-\frac{1}{2}\right)\ell\right)^{2}\phi(t)\mathrm{d}t$ (49a)
$=\sum\limits_{k=1}^{K}\int_{(k-1)L/K}^{kL/K}\left(t-\frac{k-\frac{1}{2}}{K}L\right)^{2}\phi(t)\mathrm{d}t.$ (49b)
@@
Figure 3: Convergence of achievable rate to channel capacity.
The numerical results in Fig. 4 and Fig. 5 reveal the increasing importance of gain control (realized by an AGC module in practical systems) as the resolution decreases: 1) Under high-resolution output quantization, we only require a rough estimate of the channel gain to guarantee that the loading factor remains no less than a predefined threshold, say $4$ (from the four-sigma rule of thumb [45, 47]), so that we can stay away from the waterfall; 2) Under low-resolution output quantization, such a simple strategy may increase rate loss considerably, but perfect gain control needs accurate channel estimation which is also challenging in this case.
@@
Figure 5: Impact of loading factor on achievable rate: Fixed SNR, varying resolution (the notation $\mathsf{SNR}_{q}(b)$ stands for the ‘SNR’ of a uniform quantizer with resolution $b$; see the remark after Corollary 10 for its definition, and see Table I for numerical results.)
@@ IV-B Consistency Between Rate Maximization and MSE Minimization and a GMI-MMSE Formula
From source quantization theory [48, 51] we know that, for Gaussian input, the MSE of the uniform quantizer is a strictly convex function of the step size (or loading factor) and the minimum is located at the unique solution of $\frac{\mathrm{d}\mathsf{MSE}}{\mathrm{d}\ell}=0$, denoted by $\ell^{*}$. In fact, that solution is also the unique step size that maximizes the GMI; that is, for uniform quantization at the receiver there is a consistency between achievable rate maximization and MSE minimization. This consistency further leads to a GMI-MMSE formula which connects the maximum achievable rate and the MMSE of quantization in a simple and closed form. We first give the following lemma, and then use it to establish the aforementioned findings in Theorem 8. Both results exclude the special case $b=1$ because for symmetric one-bit quantization we have $\mathcal{A}=\ell/2$, $\mathcal{B}=\pi\ell^{2}/8$, and $\gamma\equiv 1-2/\pi$ for $0<\ell<\infty$; i.e., the gain control is unnecessary since it has no impact on the achievable rate.
@@
    $0<\mathsf{mmse}=\gamma(L^{*})\leq\gamma,$ (53)
@@
First, in [48], it has been proved that the MSE is minimized if and only if the loading factor is set to be the unique solution of $\frac{\mathrm{d}\mathsf{mse}}{\mathrm{d}L}=0$, denoted by $L^{*}$. Specifically, the uniqueness of $L^{*}$ has been confirmed in [References, Sec. V-A] by showing that the second derivative of the MSE is strictly positive. Lemma 7 implies that the loading factor that satisfies $\frac{\mathrm{d}I_{\text{GMI}}}{\mathrm{d}L}=\frac{\mathrm{d}I_{\text{GMI}}}{K\mathrm{d}\ell}=0$ is also unique. Noting that the achievable rate $I_{\text{GMI}}$ is a continuous and differentiable function of $L$, from Proposition 4 we can infer that $L^{*}$ also maximizes $I_{\text{GMI}}(L)$ (it is not difficult to exclude the other possible case that $L^{*}$ minimizes $I_{\text{GMI}}(L)$). Then the first part of Theorem 8 is proved.
@@
Corollary 9: For uniform receiver quantization with resolution $b\geq 1$, the achievable rate given in (21), as a function of the loading factor $L$, satisfies
@@
where $I_{\textrm{GMI}}^{\textrm{1-bit}}$ is given in (35). Correspondingly, the parameter $\gamma$ satisfies
$0<\gamma\leq 1-\frac{2}{\pi},$ (57)
where the equality holds when $b=1$.
Proof:
The second and third equality in (56) follows directly from Proposition 4. Combining these two extreme cases with the uniqueness of the solution of $\frac{\mathrm{d}I_{\text{GMI}}}{\mathrm{d}L}=0$ when $b>1$ (see the proof of Theorem 8), we can infer that, for $0<L<\infty$, the infimum of $I_{\text{GMI}}(L)$ is $I_{\textrm{GMI}}^{\textrm{1-bit}}$. Correspondingly, $\gamma$ is upper bounded by its one-bit case, namely $1-\frac{2}{\pi}$. ∎
For the case $b=1$, from Proposition 5 it can be shown that, when $\ell=4/\sqrt{2\pi}=1.5958$, the normalized MSE achieves its minimum as $\mathsf{mmse}=1-2/\pi=\gamma$, implying that the GMI in this case can also be written as (54).
@@
The equivalence (52) established in Theorem 8 enables us to utilize existing results on MSE-optimal uniform quantization. In particular, combining (52) and results in [68, 48], we have the following corollary with respect to GMI, which implies that the optimal loading factor $L^{*}$ grows with the resolution $b$ like $2\sqrt{b\ln 2}$.
@@
For (60) and (61), we also provide new proofs, respectively, by asymptotic results on the achievable rate; see Sec. VI-A.
Remark: In quantization theory, the “SNR” for a quantizer is often defined as the ratio between the variance of the quantizer input and the MSE [42, 47]. For the optimal uniform quantizer we denote
@@
Theorem 8 implies that the maximum of the saturation rate given in (31) is
@@
$\mathsf{SNR}_{\textrm{e}}=\frac{1-\mathsf{mmse}}{\mathsf{mmse}\cdot\mathsf{SNR}+1}\mathsf{SNR}=\frac{(1-\mathsf{mmse})|h|^{2}\sigma_{x}^{2}}{\mathsf{mmse}\cdot|h|^{2}\sigma_{x}^{2}+\sigma^{2}}.$ (67)
@@
11 2 1.5958 1.5958 0.3634 0.2310 4.40 1.4604 2.11
22 4 1.9914 0.9957 0.1188 0.1155 9.25 3.0728 3.11
[…]
@@
  • SNR-limited (high resolution) scenario: $b>2\log_{10}\mathsf{SNR}$ $b_{1}$ bits and $\mathsf{SNR}_{q}\gg\mathsf{SNR}$, where $b_{1}$ can be 22~33. The irreducible loss (58) is negligible, and significant rate improvement can be obtained by increasing the SNR. In fact, we have
@@
IV-C Geometry of Optimal Uniform Quantization at the Receiver
We have shown that, in the standard transceiver architecture shown in Fig. 1, if the gain control factor $g$ is set appropriately so that the loading factor of the quantizer $q(\cdot)$ is equal to the optimal value $L^{*}$, then the achievable rate attains its maximum given by the GMI-MMSE formula (54). In fact, a geometrical interpretation for such optimal uniform quantization can be established. For simplification we first introduce an equivalent model shown in Fig. 7, where we let $\mathsf{X}=hX$, $\mathcal{E}_{x}=|h|^{2}\sigma_{x}^{2}$, $\mathsf{Y}=\sigma_{v}Y=Y\cdot\sqrt{(\mathcal{E}_{x}+\sigma^{2})/2}$, $\mathsf{V}=V^{\mathrm{R}}+\mathrm{j}V^{\mathrm{I}}$, $\mathsf{Z}=Z$, and $\mathsf{W}=\mathsf{Y}-\mathsf{V}$ (we write $\mathsf{W}(\mathsf{V})$ in Fig. 7 to emphasize that $\mathsf{W}$ is a function of $\mathsf{V}$), so that